I'm so confused, is there anyone that can help.
Solve: 5^2x-5∙5^x+6=0

Answer by lwsshak3(11628)

I'm so confused, is there anyone that can help.
Solve: 5^2x-5∙5^x+6=0
let u=5^x
u^2=5^2x
..
u^2-5u+6=0
(u-3)(u-2)=0
..
u=3=5^x
take log of both sides
xlog5=log3
x=log3/log5
x≈.6826
or
u=2=5^x
take log of both sides
xlog5=log2
x=log2/log5
x≈.4307